For a standard normal distribution, find: P(z > c) = 0.7211 (Round to two decimal places.) C=
Answer #1

Z N Normal Co,l). PCZYC) o - O,721 P( Zec) 0.7211 p(Z2C) = 1 -0.421) PCZcc) = 0.2789 using Excel C = NORM SINV (0.2789) = -0.